Given sin a = 24/25, a in quadrant 2, and cos B = -4/5, in quadrant, find A) sin(a+B) and B) cos(a+B).


1
Expert's answer
2021-04-11T13:38:14-0400

Given:

sin(a) = 24/25, a in 2 quadrant.

cos(b) = -4/5, b in 2 quadrant.





A) sin (a+b)


"\boxed{sin (a+b)=sin(a)cos(b)+cos(a)sin(b)}"

"sin (a+b)=({24\over25})*(-{4\over5})+(-{7\over25})({3\over5})"


"sin (a+b)=-{117\over125}"




B)cos(a+b)


"\boxed{cos(a+b)=cos(a)cos(b)-sin(a)sin(b)}"

"cos(a+b)=(-{7\over25})*(-{4\over5})-({24\over25})*({3\over5})"


"Cos(a+b)=-{44\over125}"